OpenStudy (anonymous):

A two digit # is 11 times its ones digits. The sum of the digits is 12. What is the number? Solve a system of equations

OpenStudy (anonymous):

two digit = #ab = 10a + b = 11b a + b = 12 a = 12 - b 10(12-b) + b = 11b 120 - 10b + b = 11b 120 = 20b b = 6 a = 6 number is 66